OpenStudy (anonymous):

Variable y varies directly with variable x, and y = 6 when x = 4. Enter the constant of variation in lowest terms. y =-----x in a fraction

OpenStudy (anonymous):

@Directrix can you help me understand?

Directrix (directrix):

6 = k * 4 so k = 6/4 = 3/2

Directrix (directrix):

So, what do you think the answer is: y =-----x @

Directrix (directrix):

@faith17smith